# Free energy of the gas of spin 1/2 fermions beyond the second order and the Stoner phase transition

### v1

@@ Abstract
In the previous work we have developed a systematic thermal (imaginary time) perturbative expansion and applying it to the relevant effective field theory computed, up to the second order in the interaction, the free energy $F$ of the diluted gas of (nonrelativistic) spin $1/2$ fermions interacting through a spin-independent repulsive two-body potential. Here we extend this computations to higher orders: assuming that the only relevant parameter specifying the interaction potential is the $s$-wave scattering length $a_{0}$, we compute the complete order $(k_{\rm F}a_{0})^{3}$ ($k_{\rm F}$ is the Fermi wave vector) contribution to the system’s free energy as a function of the numbers $N_{+}$ and $N_{-}$ of spin up and spin down fermions (i.e. as a function of its polarization) and the temperature $T$. We also extend the computation beyond a fixed order by resumming the contributions to $F$ of two infinite sets of Feynman diagrams: the so called particle-particle rings and the particle-hole rings. We find that including the second one of these two contributions has a dramatic consequence for the transition of the system from the paramagnetic to the ferromagnetic phase (the so called Stoner phase transition): in this approximation the phase transition simply disappears.
@@ 1. Introduction
In the recent years some progress has been achieved in the computation of equilibrium zero temperature properties of the gas of $N$ (nonrelativistic) fermions interacting with each other through a spin independent repulsive potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$. It was mainly related to the application to this classic [1, 2, 3, 4, 5] many-body quantum mechanics/statistical physics problem of general methods (see e.g. [6]) of the effective field theory. In this approach, initiated in the seminal paper [7] (see also [8]), the original spatially nonlocal potential $V_{\rm pot}$ assumed to be characterized by a length scale $R$ is replaced by the a priori infinite series of local interactions (written here using the standard second quantization formalism - see e.g. [9])
@@
of decreasing length dimension. The coefficients (couplings) $C_{0}$, $C_{2}$, $C_{2}^{\prime},\dots$ of the interaction (1) can be then directly determined in terms of the quantities - the scattering lengths $a_{0},a_{1},\dots$ and the effective radii $r_{0},\dots$ - parametrizing the general expansion (in powers of the relative momentum) of the amplitude of the elastic scattering of two particles. Trading the (bare) couplings of (1) for these measurable quantities which characterize the underlying potential $V_{\rm pot}$ of the binary interactions has also the effect of removing ultraviolet (UV) infinities engendered by the locality of (1). The simplifications brought in by this approach allowed first to easily reproduce [7] those terms of the perturbative expansion of the ground state energy $E$ of the system of spin $s$ fermions with equal densities of different spin projections which in the past were obtained by more traditional (and requiring considerably more work) methods of many-body quantum mechanics and to extend [10] this computation to the fourth order in the systematic, organized by the power counting rules [7], expansion in powers (in higher orders modified also by logarithms) of the dimensionless product $k_{\rm F}R$ of the system’s overall Fermi wave vector
@@
(here $n=N/V$ is the overall density of the gas of spin $s$ fermions and $g_{s}=2s+1$) and the characteristic length $R$. The same approach allowed also to compute the ground state energy of the system of spin $1/2$ fermions for different densities of fermions with the up and down spin projections [11] recovering the old result of Kanno [12] (obtained analytically for the hard core interaction potential) and to easily extend it [13] to fermions having spin $s$ greater than 1/2 (and, therefore, more possible spin projections). Finally in the works [14, 15] and [16] this computation has been extended up to the third order in the systematic expansion in powers of $k_{\rm F}R$.
@@
These results allowed to investigate more quantitatively the phase transition, called also the Stoner transition, to the ordered state in which the densities $N_{+}/V$ and $N_{-}/V$ of fermions of opposite spin projections in the case of $s=1/2$ are not equal and the system exhibits a nonzero polarization $P\equiv(N_{+}-N_{-})/N$ which, according to the standard qualitative argument based on the positivity of the repulsive interaction energy and the Pauli exclusion principle, for sufficiently strong repulsion and/or sufficiently high overall density, should be energetically (at zero temperature) favored over the state with equal densities. In this regime the system should, therefore, exhibit the property called itinerant ferromagnetism. It has been found that the clear first order character of this transition predicted by the second order of the perturbative expansion (in agreement with the general quantitative arguments given long time ago in [17]) - at sufficiently large values of the “gas parameter” $k_{\rm F}a_{0}$ two symmetric minima of $E(P)$ are formed away from $P=0$ separated from the one at $P=0$ by a finite barrier and at $(k_{\rm F}a_{0})_{\rm cr}=1.054$ they become the global minima ($|P_{\rm cr}|=0.58$) - gets appreciably weakened ($|P_{\rm cr}|$ shifts significantly towards $P=0$ and the barrier becomes much lower) if the complete order $(k_{\rm F}a_{0})^{3}$ contribution is taken into account [14, 15]. This seemed to support the results of the work [18] in which a certain class of contributions to $E$ (arising from the so-called particle-particle ring diagrams) has been resummed to all orders in $k_{\rm F}a_{0}$ finding that in this approximation the transition is continuous - the new minima start to continuously move away from $P=0$ as the gas parameter crosses some critical value $(k_{\rm F}a_{0})_{\rm cr}\approx 0.8$ ($0.858$ if only 1 hole-hole N1N-1 particle-particle parts of the $N$-th order particle-particle ring diagrams and $0.79$ if the complete $N$-th order particle-particle ring diagrams are resummed; see also [20] for a refinement of this approach). It should be added that the predictions of [18], which qualitatively are supported also by the results of a different approach [19], seem to agree quite well with the results obtained with the help of the Monte Carlo simulations [21].
@@
However if all scattering lengths $a_{\ell}$ and effective radii $r_{\ell}$ are of the same order of magnitude, $\sim\!R$, the complete (according to the power counting rules [7] which apply, strictly speaking, only to this case) order $(k_{\rm F}R)^{3}$ contribution to the ground state energy depends also on $a_{1}$ and $r_{0}$ and it has been shown in [16] that in this case the character of the phase transition (at zero temperature) predicted by this approximation depends on the relative magnitudes and signs of the parameters $a_{1}$, $r_{1}$ and $a_{0}>0$ (that is, on the more detailed characteristics of the underlying potential $V_{\rm pot}$). The computation in which to all orders resummed are only (some) contributions depending on the powers of $k_{\rm F}a_{0}$ corresponds rather to the situation encountered in physics of dilute atomic gases in which the $s$-wave scattering length $a_{0}$ (it can be of either sign) is made positive and very large compared to the remaining parameters ($|a_{0}|\gg R\sim|a_{1}|,|r_{1}|,\dots$) by exploiting properties of the Feshbach resonances (see e.g. [22]). In this case however the underlying interaction of fermions (atoms) is attractive and the scattering length $a_{0}$ is positive in the regime in which bound states composed of two fermions of opposite spins can form. The true ground state of the system is then very different than the one of nointeracting atoms (they may not be adiabatically connected to one another in the thermodynamic limit implicit in the field theory approach) which is used in perturbative computations performed within the effective field theory approach. Although the formulae for the system’s energy density $E/V$ obtained perturbatively (or with the help of a resummation) by expanding around the ground state of noninteracting fermions can seem to imply the transition to the ordered state, they apply at best to a metastable (from the thermodynamic point of view) state of the system and in real experiments the transition (which in principle could occur in a metastable state [23]) to the ferromagnetic state has in fact not been observed [24, 25, 26] due to the too rapid formation (at very low temperatures at which these experiments were carried out) of atomic dimers (bosons). It can be also remarked that in the situation in which the underlying interaction is attractive (despite giving rise to a positive $s$-wave scattering length) the mentioned qualitative argument for the occurrence of the transition no longer applies and the expectation that the transition should occur is mainly based on the textbook mean field correction to the energy density [4, 5] (equivalent to the first order correction of the perturbative expansion) which depends only on the $s$-wave scattering length $a_{0}$.
@@
Computation of the ground state energy density $E/V$ as a function of the densities of fermions with different spin projections allows only to investigate the equilibrium properties of the system of interacting fermions at zero (or very low) temperatures. It is however of interest to determine its behavior also at nonzero temperatures. (In the context of the physics of atomic gases it is physically clear that formation of atomic dimers, which at very low temperatures makes observation of the phase transition to the ordered phase impossible, at higher temperatures should be less important.) This requires computing one of the thermodynamic potentials of the system of interacting fermions. So far such a computation of the free energy $F$ has been done [27] only up to the second order, i.e. up to terms of order $(k_{\rm F}a_{0})^{2}$, using the old-fashioned thermal perturbation theory (see e.g. [28], par. 32) based on the ordinary second order Rayleigh - Schrödinger perturbative expression for energy levels entering the statistical sum. In [29] we have recovered this second order expression for $F$ using the systematic thermal perturbative expansion (exploiting the imaginary time formalism [9]) and reproduced the thermal characteristics of the Stoner phase transition it predicts (pointing out however problems - not discussed in [27] - with accurate numerical determination of the critical values of the polarization), but have encountered a technical problem which prevented us to immediately extend the computation to higher orders. Here we show how this problem can be resolved and applying the developed systematic thermal expansion to the first term of the effective field theory interaction (1) we derive the formulae allowing to compute numerically the complete order $(k_{\rm F}a_{0})^{3}$ corrections to the free energy $F(T,V,N_{+},N_{-})$. With additional work, including the contributions of the next two terms of (1) it would, of course, be possible to compute the complete order $(k_{\rm F}R)^{3}$ correction to the free energy.
@@
Instead of completing the order $(k_{\rm F}R)^{3}$ corrections to the free energy which would allow to investigate in more details the thermal profile of the Stoner phase transition to the ordered state induced by truly repulsive spin independent two-body potentials $V_{\rm pot}$ which necessarily give rise to the parameters $a_{1}$ and $r_{0}$ of comparable magnitude to that of $a_{0}$ (all $\sim R)$, we in this paper profit from the possibility provided by the simpler structure of the terms of the expansion generated by the imaginary time formalism and resum to all orders in $k_{\rm F}a_{0}$ not only the contributions to the temperature-dependent free energy $F$ of the particle-particle ring diagrams (done for zero temperature in the papers [18, 20, 30]) but also of the particle-hole ring diagrams.
@@
Including in the free energy $F(T,V,N,P)$ only the resummed contribution of the particle-particle ring diagrams we recover for $T=0$ the results of the works [18, 30] and can show how they are modified at nonzero temperatures not exceeding the Fermi temperature (we expect that the results obtained within the effective field theory should be valid for temperatures in this range). However as we show, inclusion of the contribution of the resummed particle-hole diagrams changes the situation drastically: the phase transition to the ordered state simply disappears (the minimum of $F$ is at $P=0$ for all values of the parameter $k_{\rm F}a_{0}$ and all temperatures). This is a somewhat surprising result and we comment on its possible meaning in the Conclusions.
@@ 2. The formalism
The natural statistical formalism in which to compute equilibrium properties of the gas of fermions the interaction of which preserve their spins and therefore the numbers $N_{\pm}$ of spin up and spin down particles, is the Grand Canonical Ensemble with two independent chemical potentials $\mu_{\pm}$. The relevant statistical operator is then
@@
with $\varepsilon_{\mathbf{p}}=\hbar^{2}{\mathbf{p}}^{2}/2m_{f}$. Standard systematic thermodynamical perturbative expansion [9, 29] gives the potential $\Omega$ in the form of the series in powers of the interaction $\hat{V}_{\rm int}$
@@
Here $\hat{V}_{\rm int}^{I}(\tau)=e^{\tau\hat{K}_{0}}\hat{V}_{\rm int}e^{-\tau\hat{K}_{0}}$ is the interaction operator in the (imaginary time) interaction picture, ${\rm T}_{\tau}$ is the chronological ordering and $\hat{\rho}^{(0)}$ is the statistical operator of the noninteracting system. The superscript “con” means that only connected contributions (Feynman diagrams) should be taken into account. The first term in (6) is the textbook [4, 5] grand thermodynamical potential of the nointeracting system
@@
Owing to the thermal analog of the Wick formula (see e.g. [9]) computation of the successive terms $\Omega^{(N)}$ of the expansion (6) reduces to drawing all possible connected Feynman diagrams with $N$ interaction vertices arising from $\hat{V}_{\rm int}$ joined by oriented lines and integrating over positions $\mathbf{x}$} and “times” $\tau$ labeling these vertices the corresponding products of free the thermal propagators
@@
associated with (oriented) lines connecting vertices of the diagram. The resulting “momentum” space Feynman rules are almost identical with the ordinary ones except that the integrations over frequencies (energies) are replaced by summations over the (fermionic) Matsubara frequencies $\omega_{n}^{\rm F}=(\pi/\beta)(2n+1)$, $n\in{\mathbb{Z}}$.
@@
Higher order contributions to the potential $\Omega$ can also be systematically computed. If $\hat{V}_{\rm int}$ in (4) were the true, spatially nonlocal, two-body interaction (corresponding to a two potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$, where ${\mathbf{x}}_{i}$ are the positions of fermions), the successive terms of the expansion (6) would be (ultraviolet) finite. If $\hat{V}_{\rm int}$ is the local interaction (1) of the effective theory, the successive terms of the expansion (6) involve ultraviolet divergences and have to be regularized. As in our previous works we will employ for this purpose the cutoff $\Lambda$ on the wave vectors of virtual particles. Finite (in the limit $\Lambda\rightarrow\infty$) contributions to the potential $\Omega$ (a physical quantity) are then obtained by systematically expressing the (bare) couplings of $\hat{V}_{\rm int}$ in terms of other measurable (physical) quantities. As it is customary, and in line (at least when the gas is very diluted) with the physical intuition that properties of the gas are mainly determined by elastic two-body collisions of its constituents, one expresses the couplings $C_{0}$, $C_{2}$, etc. of (1) in terms of the measurable quantities related to such a scattering process, namely in terms of the scattering lengths $a_{0}$, $a_{1}$, effective ranges $r_{0}$, etc. [7]. In this work we will only need to express the coupling $C_{0}$ in this way; the relevant formula obtained by matching the amplitude of the elastic fermion-fermion scattering computed perturbatively using the first term of the interaction (1) onto the general form of the same amplitude parameterized by $a_{0}$, $a_{1},\dots$ and $r_{0},\dots,$ reads [11, 14, 10]
@@
In principle, in each order of the expansion (6) to construct the free energy one should invert the relations (13) to obtain the chemical potentials as functions of the particle numbers $N_{+}$ and $N_{-}$ and insert these in the formula $F=\Omega+\mu_{+}N_{+}+\mu_{-}N_{-}$. Thus the values of the chemical potentials $\mu_{+}$ and $\mu_{-}$ change with each successive order of the expansion and the procedure of constructing the free energy looks rather cumbersome. It turns out, however, that in the systematic expansion this procedure simplifies considerably: it amounts in effect to using the chemical potentials $\mu_{\pm}^{(0)}$ determined by inverting the formula (13) with $\Omega$ replaced by $\Omega^{(0)}$ given by (7) and omitting in the expansion (6) those diagrams which give vanishing contribution in computing the corrections $\Delta E$ to the ground state energy $E=E^{(0)}+\Delta E$ of the system of interacting particles using the ordinary Dyson expansion of the formula [7, 10, 11, 14, 15] ($T$ stands here for time, not for the temperature)
@@
can be also easily taken into account in this formalism by including it in the free Hamiltonian $\hat{H}_{0}$ which amounts to shifting the chemical potentials $\mu_{\pm}\rightarrow\tilde{\mu}_{\pm}\equiv\mu_{\pm}\pm{\cal H}$ in $\hat{K}_{0}$ given by (4). The free energy is then given as the series

### v2

# Free energy of the gas of spin-1/2 fermions beyond the second order and the Stoner phase transition

@@ Abstract
Applying the previously developed systematic thermal (imaginary time) perturbative expansion to the relevant effective field theory we compute the free energy $F$ of the diluted gas of (nonrelativistic) spin $1/2$ fermions interacting through a spin-independent repulsive two-body potential as a function of the numbers $N_{+}$ and $N_{-}$ of spin up and spin down fermions (i.e. as a function of the system’s polarization) and the temperature $T$. We give the complete order $(k_{\rm F}a_{0})^{3}$ ($k_{\rm F}$ is the Fermi wave vector and $a_{0}$ is the $s$-wave scattering length characterizing the interaction potential) contribution to $F$. We also extend the computation beyond a fixed order by resumming to all orders in the parameter $k_{\rm F}a_{0}$ the contributions to $F$ of two infinite sets of Feynman diagrams: the so-called particle-particle rings and the particle-hole rings. We find that including the second one of these two contributions has a dramatic consequence for the transition of the system from the paramagnetic to the ferromagnetic phase (the so called Stoner phase transition): in this approximation the phase transition simply disappears. This result does not contradict the expectation that a transition to the magnetically ordered state should occur in truly repulsive systems. The $p$-wave and higher scattering lengths, as well as other parameters chacterizing the interaction potential, are in such systems generally of the same order of magnitude as $a_{0}$ and contributions depending on them should be, therefore, also included in $F$. Our results may, however, have implications for the search of the itinerant ferromagnetism of cold atomic gases in which large $a_{0}$, much larger than all other parameters, is artificially created by exploiting the physics of the Feshbach resonance.
@@ 1. Introduction
In recent years some progress has been achieved in the computation of equilibrium zero temperature properties of the gas of $N$ (nonrelativistic) fermions interacting with each other through a spin independent repulsive potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$. It was mainly related to the application to this classic [1, 2, 3, 4, 5, 6] many-body quantum mechanics and statistical physics problem of general methods (see e.g. [7]) of the effective field theory. In this approach, initiated in the seminal paper [8] (see also [9]), the original spatially nonlocal potential $V_{\rm pot}$ assumed to be characterized by a length scale $R$ is replaced by the a priori infinite series of local interactions (written here using the standard second quantization formalism - see e.g. Ref. [6] - with $\hat{\psi}_{\pm}$ being the field operators of spin up and down fermions)
@@
of decreasing length dimension. The coefficients (couplings) $C_{0}$, $C_{2}$, $C_{2}^{\prime},\dots$ of the interaction (1) can be then directly determined in terms of the quantities - the scattering lengths $a_{0},a_{1},\dots$ and the effective radii $r_{0},\dots$ - parametrizing the general expansion (in powers of the relative momentum) of the amplitude of the elastic scattering of two particles. Trading the (bare) couplings of (1) for these measurable quantities which characterize the underlying potential $V_{\rm pot}$ of the binary interactions also has the effect of removing UV infinities engendered by the locality of (1). The simplifications brought in by this approach allowed first to easily reproduce [8] those terms of the perturbative expansion of the (ground-state) energy density $E/V$ of the system of spin $s$ fermions with equal densities of different spin projections which in the past were obtained by more traditional (and requiring considerably more work) methods of many-body quantum mechanics (such as the ones based on Green’s functions or on the Galitskii approach) [6] and to extend [10] this computation to the fourth order in the systematic, organized by the power counting rules [8], expansion in powers (in higher orders modified also by logarithms) of the dimensionless product $k_{\rm F}R$ of the system’s overall Fermi wave vector
@@
(here $n=N/V$ is the overall density of the gas of spin $s$ fermions and $g_{s}=2s+1$) and the characteristic length $R$. The same approach allowed also to compute up to order $(k_{\rm F}a_{0})^{2}$ the ground state energy of the system of spin $1/2$ fermions for different densities of fermions with the up and down spin projections [11] recovering the old result of Kanno [12] (obtained analytically for the hard core interaction potential) and to easily extend it [13] to fermions having spin $s$ greater than 1/2 (and, thus, more possible spin projections). Finally, in Refs. [14, 15] and slightly later in [16] this computation has been extended up to the third order in the systematic expansion in powers of $k_{\rm F}R$.
@@
These technical achievements allowed to investigate more quantitatively the phase transition, also called the Stoner transition, to the ordered state in which the densities $N_{+}/V$ and $N_{-}/V$ of fermions of opposite spin projections in the case of $s=1/2$ are not equal and the system exhibits a nonzero polarization $P\equiv(N_{+}-N_{-})/N$. According to the standard qualitative argument based on the positivity of the repulsive interaction energy and the Pauli exclusion principle, for sufficiently strong repulsion and/or sufficiently high overall density, such an ordered state should be energetically (at zero temperature) favored over the state with equal densities. In this regime the system should, therefore, exhibit the property called itinerant ferromagnetism. It has been found that the clear first order character of this transition predicted by the second order of the perturbative expansion (in agreement with the general quantitative arguments given long ago in Ref. [17]) - at sufficiently large values of the “gas parameter” $k_{\rm F}a_{0}$ two symmetric minima of $E(P)$ are formed away from $P=0$ separated from the one at $P=0$ by a finite barrier and at $(k_{\rm F}a_{0})_{\rm cr}=1.054$ they become the global minima ($|P_{\rm cr}|=0.58$) - gets significantly weakened ($|P_{\rm cr}|$ shifts significantly toward $P=0$ and the barrier becomes much lower) if the complete order $(k_{\rm F}a_{0})^{3}$ contribution is taken into account [14, 15]. This seemed to be in line with the results of the work [18] in which a certain class of contributions to $E/V$ (arising from the so-called particle-particle ring diagrams) has been resummed to all orders in $k_{\rm F}a_{0}$ finding that in this approximation the transition is continuous - the new minima start to continuously move away from $P=0$ as the gas parameter crosses some critical value $(k_{\rm F}a_{0})_{\rm cr}\approx 0.8$ ($0.858$ if only 1 hole-hole N1N-1 particle-particle parts of the $N$-th order particle-particle ring diagrams and $0.79$ if the complete $N$-th order particle-particle ring diagrams are resummed; see also Ref. [19] for a refinement of this approach). It should be added that the predictions of Ref. [18], which qualitatively seem to be supported by the results of a different approach [20], agree at first sight quite well with the results obtained with the help of the Monte Carlo simulations [21].
@@
However if all scattering lengths $a_{\ell}$ and effective radii $r_{\ell}$ are of the same order of magnitude $\sim\!R$, the complete (according to the power counting rules [8] which apply, strictly speaking, only to such a case) order $(k_{\rm F}R)^{3}$ contribution to the energy density depends also on $a_{1}$ and $r_{0}$ and it has been shown in Ref. [16] that in this case the character of the phase transition (at zero temperature) predicted by this approximation depends on the relative magnitudes and signs of the parameters $a_{1}$, $r_{1}$ and $a_{0}>0$ (that is, on the more detailed characteristics of the underlying potential $V_{\rm pot}$). The computation (such as the one done in Ref. [18]) in which to all orders resummed are (some) contributions depending only on powers of $k_{\rm F}a_{0}$ correspond rather to the situation encountered in physics of dilute atomic gases in which the $s$-wave scattering length $a_{0}$ (it can be of either sign) is made positive and very large compared to the remaining parameters ($|a_{0}|\gg R\sim|a_{1}|,|r_{1}|,\dots$) by exploiting properties of the Feshbach resonances (see e.g. Ref. [22]). In this case however the underlying interaction of fermions (atoms) is attractive and the scattering length $a_{0}$ is positive in the regime in which bound states composed of two fermions of opposite spins can form. The true ground state of the system is then very different from the one of nointeracting atoms (they may not be adiabatically connected to one another in the thermodynamic limit implicit in the field theory approach) which is used in perturbative computations performed within the effective field theory approach. In view of that the energy density $E/V$ obtained perturbatively by expanding around the ground-state of the system of noninteracting fermions most probably is the energy density of the metastable (from the thermodynamic point of view) state in which a real gas of cold fermionic atoms finds itself just after switching the external magnetic field so that the $s$-wave scattering length $a_{0}$ becomes positive and large. However, although the dependence on the polarization of this energy density indicates that the transition to the ferromegnetic state (which in principle could occur in a metastable state [23]) should occur, a too short lifetime of the gas metastable state may prevent its observation in real systems. Therefore the failure of the attempts to observe such a transition [24, 25, 26] in experiments is usually attributed to the too rapid formation (at very low temperatures at which these experiments were carried out) of atomic dimers (bosons). It should however be remarked that in the situation in which the underlying interaction is attractive (despite giving rise to a positive $s$-wave scattering length) the mentioned qualitative argument for the occurrence of the transition no longer applies and the expectation that the transition should occur is mainly based on the textbook mean field correction to the energy density [4, 5] (equivalent to the first order correction of the perturbative expansion) which depends only linearly on the $s$-wave scattering length $a_{0}$.
@@
Computation of the ground-state energy density $E/V$ as a function of the densities of fermions with different spin projections allows only to investigate the equilibrium properties of the system of interacting fermions at zero (or a very low) temperature. It is however of interest to determine its behavior also at nonzero temperatures. (In the context of the physics of atomic gases it is physically clear that at higher temperatures the mentioned formation of atomic dimers should be less important.) This requires computing one of the thermodynamic potentials of the system of interacting fermions. So far such a computation of the free energy $F$ has been done [27] only up to the second order, i.e. up to terms of order $(k_{\rm F}a_{0})^{2}$, using the old-fashioned thermal perturbation theory (see e.g. Ref. [28], paragraph 32) based on the ordinary second order Rayleigh - Schrödinger perturbative expression for energy levels entering the statistical sum. In Ref. [29] we have applied to the effective field theory (1) the standard thermal perturbative expansion (exploiting the imaginary time formalism - see e.g. Ref. [6]) which is routinely used to compute thermodynamic characteristics of much more complex systems like e.g. the quark-gluon plasma. Using it we recovered this second order expression for $F$ and reproduced the thermal characteristics of the Stoner phase transition it predicts (however, pointing out problems - not discussed in [27] - with accurate numerical determination of the critical values of the polarization), but have encountered a technical problem which prevented us from immediately extending the computation to higher orders. Here we show how this problem can be resolved and applying the developed systematic thermal expansion to the first term of the effective field theory interaction (1) we derive the formulas allowing to compute numerically the complete order $(k_{\rm F}a_{0})^{3}$ corrections to the free energy $F(T,V,N_{+},N_{-})$. With additional work, including the contributions of the next two terms of (1) it would be possible to compute the complete order $(k_{\rm F}R)^{3}$ correction to the free energy.
@@
In this paper, instead of completing the order $(k_{\rm F}R)^{3}$ correction to the free energy (which would allow to investigate in more detail the thermal profile of the Stoner phase transition to the ordered state induced by truly repulsive spin-independent two-body potetials $V_{\rm pot}$ which necessarily give rise to the parameters $a_{1}$ and $r_{0}$ of comparable magnitude to that of $a_{0}$), we profit from the possibility provided by the simpler structure of the terms of the expansion generated by the imaginary time formalism and resum to all orders in $k_{\rm F}a_{0}$ not only the contributions to the temperature-dependent free energy $F$ of the particle-particle ring diagrams (done for zero temperature in Refs. [18, 19, 30]) but also of the particle-hole ring diagrams. The thermodynamic potential obtained in this way is most probably the free energy of the system of fermionic atoms in a metastable state on the BEC side of the Feshbach resonance and can therefore serve to study its properties.
@@
Including in the free energy $F(T,V,N,P)$ only the resummed contribution of the particle-particle ring diagrams we recover for $T=0$ the results of the works [18, 30] and can study how they are modified at nonzero temperatures not exceeding the Fermi temperature (we expect that the results obtained within the effective field theory should be valid for temperatures in this range). However we find that the inclusion of the contribution of the resummed particle-hole diagrams changes the situation drastically: the phase transition to the ordered state simply disappears (the minimum of $F$ is at $P=0$ for all values of the parameter $k_{\rm F}a_{0}$ and all temperatures). This is a somewhat surprising result and we comment on its possible meaning for the problem of itinerant ferromagnetism in system of cold gases in the Conclusions.
@@ 2. The formalism
The natural statistical formalism in which to compute equilibrium properties of the gas of fermions the interaction of which preserves their spins and therefore the numbers $N_{\pm}$ of spin up and spin down particles, is the grand canonical ensemble with two independent chemical potentials $\mu_{\pm}$. The relevant statistical operator is then
@@
with $\varepsilon_{\mathbf{p}}=\hbar^{2}{\mathbf{p}}^{2}/2m_{f}$. The standard systematic thermodynamical perturbative expansion [6, 29] gives the potential $\Omega$ in the form of the series in powers of the interaction $\hat{V}_{\rm int}$
@@
Here $\hat{V}_{\rm int}^{I}(\tau)=e^{\tau\hat{K}_{0}}\hat{V}_{\rm int}e^{-\tau\hat{K}_{0}}$ is the interaction operator in the (imaginary time) interaction picture, ${\rm T}_{\tau}$ is the chronological ordering and $\hat{\rho}^{(0)}$ is the statistical operator of the noninteracting system. The superscript “con” means that only connected contributions (Feynman diagrams) should be taken into account. The first term in (6) is the textbook [4, 5] grand thermodynamical potential of the system of nointeracting fermions
@@
Owing to the thermal analog of the Wick formula (see e.g. Ref. [6]) computation of the successive terms $\Omega^{(N)}$ with $N\geq 1$ of the expansion (6) reduces to drawing all possible connected Feynman diagrams with $N$ interaction vertices arising from $\hat{V}_{\rm int}$ joined by oriented lines and integrating over positions $\mathbf{x}$} and “times” $\tau$ labeling these vertices the corresponding products of the free thermal propagators
@@
associated with (oriented) lines connecting vertices of each diagram. The resulting “momentum” space Feynman rules are almost identical with the ordinary ones except that integrations over frequencies (energies) are replaced by summations over the (fermionic) Matsubara frequencies $\omega_{n}^{\rm F}=(\pi/\beta)(2n+1)$, $n\in{\mathbb{Z}}$.
@@
Higher-order contributions to the potential $\Omega$ can also be systematically computed. If $\hat{V}_{\rm int}$ in (4) were the true, spatially nonlocal, two-body interaction [corresponding to a two-body potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$, where ${\mathbf{x}}_{i}$ are the positions of fermions], the successive terms of the expansion (6) would be (ultraviolet) finite. If $\hat{V}_{\rm int}$ is the local interaction (1) of the effective theory, the successive terms of the expansion (6) involve ultraviolet divergences and have to be regularized. As in our previous works we will employ for this purpose the cutoff $\Lambda$ on the wave vectors of virtual particles. Finite (in the limit $\Lambda\rightarrow\infty$) contributions to the potential $\Omega$ (a physical quantity) are then obtained by systematically expressing the (bare) couplings of $\hat{V}_{\rm int}$ in terms of other measurable (physical) quantities. As it is customary, and in line (at least when the gas is very diluted) with the physical intuition that properties of the gas are mainly determined by elastic two-body collisions of its constituents, one expresses the couplings $C_{0}$, $C_{2}$, etc. of (1) in terms of the measurable quantities related to such a scattering process, namely in terms of the scattering lengths $a_{0}$, $a_{1}$, effective ranges $r_{0}$, etc. [8]. In this work we will only need to express the coupling $C_{0}$ in this way; the relevant formula obtained by matching the amplitude of the elastic fermion-fermion scattering computed perturbatively using the first term of the interaction (1) onto the general form of the same amplitude parameterized by $a_{0}$, $a_{1},\dots$ and $r_{0},\dots,$ reads [11, 14, 10]
@@
In principle, in each order of the expansion (6) to construct the free energy one should invert the relations (13) to obtain the chemical potentials as functions of the particle numbers $N_{+}$ and $N_{-}$ and insert these in the formula $F=\Omega+\mu_{+}N_{+}+\mu_{-}N_{-}$. Thus the values of the chemical potentials $\mu_{+}$ and $\mu_{-}$ change with each successive order of the expansion and the procedure of constructing the free energy looks at first sight rather cumbersome. It turns out, however, that in the systematic expansion this procedure simplifies considerably: it amounts in effect to using the chemical potentials $\mu_{\pm}^{(0)}$ determined by inverting the formula (13) with $\Omega$ replaced by $\Omega^{(0)}$ given by (7) and omitting in the expansion (6) those diagrams which give vanishing contribution in computing the corrections $\Delta E$ to the ground state energy $E=E^{(0)}+\Delta E$ of the system of interacting particles using the ordinary Dyson expansion of the formula [8, 10, 11, 14, 15] ($T$ stands here for time, not for the temperature, and ${\rm T}_{t}$ means the ordinary chronological ordering)
@@
can also be easily taken into account in this formalism by including it in the free Hamiltonian $\hat{H}_{0}$ which amounts to shifting the chemical potentials $\mu_{\pm}\rightarrow\tilde{\mu}_{\pm}\equiv\mu_{\pm}\pm{\cal H}$ in $\hat{K}_{0}$ given by (4). The free energy is then given as the series